2. Consider the economy with only one bank. The table below gives the information of the economy. Reserve ratio is 5% and constant.

Amount of Stocks 8000
Amount of Bonds 7000
Amount of Currency 2000
Reserves of the Bank 200

2.a. Answer how much the deposits in the bank is and how much the money supply in the economy is.

              Deposits:

              Money supply
Answer #1


Question 2

(a)

Reserve ratio = 5% or 0.05

Reserve of banks = 200

Reserve ratio = Reserve/Deposits

0.05 = 200/Deposits

Deposits = 200/0.05

Deposits = 4,000

Thus,

The deposits is 4,000.

Calculate the money supply -

Money supply = Amount of currency + Deposits

Money supply = 2,000 + 4,000 = 6,000

Thus,

The money supply is 6,000.